# Automated Counting of Spanning Trees for Several Infinite families of Graphs

### v2

@@ Abstract
Using the theoretical basis developed by Yao and Zeilberger, we consider certain graph families whose structure results in a rational generating function for sequences related to spanning tree enumeration. Said families are Powers of Cycles and Powers of Path; later, we briefly discuss Torus graphs and Grid graphs. In each case we know, a priori, that the set of spanning trees of the family of graphs can be described in terms of a finite-state-machine, and hence there is a finite transfer-matrix that guarantees the generating function is rational. Finding this “grammar”, and hence the transfer-matrix is very tedious, so a much more efficient approach is to use experimental mathematics. Since computing numerical determinants is so fast, one can use the matrix tree theorem to generate sufficiently many terms, then fit the data to a rational function. The whole procedure can be done rigorously a posteriori.
@@
1 Introduction and Background
A subgraph $T$ of a graph $G$ such that $T$ is a tree with $V(T)=V(G)$ is called a spanning tree of $G$. If $G$ is connected, simple, and not a tree, then it contain several spanning trees. When we count the number of spanning trees of a graph in this paper, we consider two isomorphic trees with different vertex labels to be different trees; for example, the complete graph on three vertices $K_{3}$ has three spanning trees, not one. For a connected graph $G$, we denote its number of spanning trees of $G$ by $\tau(G)$. Cayley’s formula counts the number of spanning trees of a complete graph on $n$ vertices to be $n^{n-2}=\tau(K_{n})$.
The generating function $f(x)$ for a sequence $(a_{0},a_{1},\ldots)$ is the formal power series obtained by setting the sequence terms as its coefficients:
@@ 1 Introduction and Background
Whenever a sequence has a recurrence of finite length, with constant coefficients, it is called C-finite. In other words, $(a_{0},a_{1},\ldots)$ is C-finite (of order $r$) if there is a fixed $r$ and coefficients $c_{0},\ldots,c_{r-1}$ with $c_{0}\neq 0$ such that
$a_{n+r}=c_{r-1}a_{n+r-1}+\ldots+c_{0}a_{n}$
@@
[2][Kauers-Paule, Thm. 4.3] A sequence $(a_{0},a_{1},\ldots)$ is C-finite (of order $r$) with recurrence
$a_{n+r}=c_{r-1}a_{n+r-1}+\ldots+c_{0}a_{n}$
@@
$\sum a_{n}x^{n}=\frac{p(x)}{1-c_{r-1}x+\ldots-c_{1}x^{r-1}-c_{0}x^{r}}$
for some polynomial $p(x)$ of degree at most $r-1$.
In fact, the polynomial $p(x)$ in the previous Theorem is determined by the initial values $(a_{0},\ldots,a_{r-1})$ of the sequence. Later, we state the starting index of (the underlying graph in) our sequence to disambiguate $p(x)$ in the sequence’s rational generating function.
Due to the recursive nature of the graph families we consider, we presume that there is a finite Transfer Matrix which describes the corresponding sequence (see [4] for more details). Thanks to this, we can find the sequence’s rational generating function by computing a large number of terms in the sequence. The guessing is done in a Maple procedure, described in the same paper [4], called GuessRec.
To generate the terms of whichever sequence we consider, we apply Kirchhoff’s Matrix Tree Theorem, which allows us to compute the number of spanning trees of a(n $n$-vertex) graph by looking at its Laplacian matrix, taking an $n-1$ by $n-1$ (matrix) minor, and computing its determinant. Below, we define the Laplacian matrix of a graph for the reader’s convenience.
If $G$ is a (simple) graph with vertices $v_{1},\ldots,v_{n}$, its Laplacian matrix is a symmetric matrix with its entries defined by
@@
Afterwards, we describe similar methods that we used to count the total number of leaves (across all spanning trees) for members in the same families (in Section 3). In addition, we consider an asymptotic constant relating total leaves to total spanning trees in any families (for which the constant is well-defined) in Section 4. We state our main experimental results in Section 5 and Section 6, then briefly consider an additional pair of graph families in Section 7. In Section 8 and Section 9, we include experimental verification of our results’ correctness and then a description of the accompanying Maple package that we used. Additional outputs and code are available at [1]. Finally, we conclude with some conjectures in Section 10
@@ 2 Preliminaries
For ease of notation, we assume all graphs are simple. We write $[n]:=\{1,\ldots,n\}$.
@@
From the definition, we see that if $d$ is the diameter of a connected graph $G$ (defined as $d:=\max_{u,v\in V(G)}|u,v|$), then $G^{d}$ is the complete graph on $|V(G)|$ vertices. In some cases, it’s natural to exclude complete graphs and hence consider only $G^{k}$ when $k<d$. Additionally, one may note that if there are two shortest paths (in $G$) between two vertices $u,v$ of length $m>1$, then $G^{m}$ may be better viewed as a multigraph where there are two edges between $u,v$.
$\mathcal{G}_{r}$ :=$:=\{{C_{n}}^{r}\ :\ r<\text{diam }C_{n}\}$
$\mathcal{H}_{r}$ :=$:=\{{P_{n}}^{r}\ :\ r<\text{diam }P_{n}\}.$
The condition in each construction ensures that our graph families do not contain any "unnatural" complete graphs. For those cases, we can once again recall Cayley’s Formula: $\tau(K_{n})=n^{n-2}$. Since the diameter of a cycle (resp. path) increases with the number of vertices, the condition $r<\text{diam }C_{n}=\left\lfloor n/2\right\rfloor$ is implicitly a lower bound on the number of vertices. Explicitly,
$\mathcal{G}_{r}$ =$=\{{C_{n}}^{r}\ :\ n\geq 2r+1\}$
$\mathcal{H}_{r}$ =$=\{{P_{n}}^{r}\ :\ n\geq r+2\}.$
As discussed in the introduction, the initial values of a C-finite sequence determine the numerator of its rational generating function. The minimum values of $n$ above will mark the beginning of that family’s sequence.
Since paths and cycles are ubiquitous in graph theory, they are the focus of this paper. We provide a quick application below.
Let $G$ be a connected graph. Suppose that we wish to estimate the number of spanning trees of its $k$-th power graph, $\tau(G^{k})$. One way to estimate $\tau(G^{k})$ is to fix a vertex $v\in V(G)$, let vertex sets $H_{1},\ldots,H_{N}$ be the components of $G-v$ and $m_{i}(v):=\max_{u\in H_{i}}|u,v|$. Then, deduce the bound:
@@ 3 Counting Total Number of Leaves
The next proposition plays a key role in our implementation for computing a parameter for graph families that we call the B-Z constant. The idea is to count the total number of leaves (across all spanning trees) in a graph by counting how many times each vertex appears as a leaf. For each vertex in $G$, remove the vertex from $G$ and count the number of spanning trees in the resulting graph.
@@
We say $G$ is vertex-transitive if for any $u,v\in V(G)$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u)=v$ and $\varphi(v)=u$.
@@ 4 The B-Z Constant: Average Number of Leaves per Vertex
whenever the limit exists. This is the limit of the average number of leaves in a random spanning tree (and normalized by number of vertices) in the discussed family. For the infinite family of complete graphs $K_{n}$, the limit is famously equal to $1/e$ (see below). Since the bound $\frac{\sum|\mathcal{L}(T)|}{n\tau(G)}\leq 1$ holds for all $G$, we see that the B-Z constant only fails to exist for those graph families with members whose underlying spanning trees are radically different. The rest of the section provides a few examples of B-Z constants and our approach for computing them.
The B-Z constants for both the family of Cycles and the family of Paths is equal to 00, since every tree (a path) has a constant number of leaves (two). A star graph is a connected graph where all edges share the same vertex. When $\mathcal{G}$ is the family of star graphs, its B-Z constant is $1$ because the unique tree of a star graph (itself) has $n-1$ leaves.
Thanks to Corollary 3.0.2, computation of the B-Z constant for vertex-transitive graphs is made easier. Our implementation uses this optimization when appropriate. Incidentally, it’s easy to compute the B-Z constant for complete graphs by using Corollary 3.0.2 in conjunction with Cayley’s Formula:
@@
  • Proof.
    Recall Cayley’s formula, which states $\tau(K_{n})=n^{n-2}$ for $n\geq 2$. With $\mathcal{T}:=\mathcal{T}(K_{n})$, Corollary 3.0.2 tells us that $\sum|\mathcal{L}(T)|=n\cdot(n-1)^{n-2}$. Hence,
@@
Earlier, we noted that star graphs have B-Z constant equal to $1$. The next proposition observes that we can find a graph family with B-Z constant equal to any rational number (in the interval $[0,1]$) by slightly modifying star graphs.
@@
    To construct each member $G_{k}\in\mathcal{G}$ (assuming $k\geq 1$): begin with a star graph $S_{pk}$ with $pk$ leaves; then, subdivide $(q-p)k$ edges in $S_{pk}$ and call this $G_{k}$. Note that $|V(G_{k})|=qk+1$ and that $G_{k}$ has $pk$ leaves. Furthermore, $G_{k}$ is a tree so it has exactly one spanning tree. It follows that the B-Z constant for $\mathcal{G}$ is $\lim_{k\to\infty}\frac{pk}{qk+1}=\frac{p}{q}$.
In upcoming sections, we state the B-Z constants of several graph families. It’s possible, as we did in later sections, to compute the B-Z constant of a graph family from its asymptotic behavior in the total number of leaves and number of spanning trees (which can be obtained from their respective generating functions). We used standard residue calculations, implemented in procedure BZc from our Maple package (Section 9).
@@ 5.1 Generating Functions for the Number of Spanning Trees
In this section, we include our results for the Number of Spanning Trees for the family $\mathcal{G}_{r}$, with $2\leq r\leq 5$.
Theorem 5.1.1.
The generating function $f(t)$ for the number of spanning trees in $\mathcal{G}_{2}$ is
$\frac{-36t^{5}+132t^{4}+46t^{3}-353t^{2}-116t+125}{(t+1)^{2}(t^{2}-3t+1)^{2}}$
Theorem 5.1.2.
The generating function $f(t)$ for the number of spanning trees in $\mathcal{G}_{3}$ is
$\frac{N_{3}}{(t-1)^{2}(t^{4}+3t^{3}+6t^{2}+3t+1)^{2}(t^{4}-4t^{3}-t^{2}-4t+1)^{2}}$
where
$N_{3}$ $:=-3072t^{17}+11683t^{16}+26868t^{15}+60636t^{14}-356682t^{13}-844329t^{12}-1651344t^{11}-104646t^{10}+813834t^{9}+$
$16807\displaystyle 3128248t^{8}+1452330t^{7}+512250t^{6}-1392528t^{5}-1049445t^{4}-579514t^{3}-54068t^{2}+15716t+$ 16807 .
Theorem 5.1.3.
The generating function $f(t)$ for the number of spanning trees in $\mathcal{G}_{4}$ is
$\frac{N_{4}}{D_{4}}$
where $N_{4}$ is a polynomial of degree $53$ and
$D_{4}$ $=(t+1)^{2}(t^{6}-3t^{5}+6t^{4}-10t^{3}+6t^{2}-3t+1)^{2}(t^{8}-4t^{7}-17t^{6}+8t^{5}+49t^{4}+8t^{3}-17t^{2}-4t+1)^{2}$
$(t^{12}+3t^{11}+12t^{10}+28t^{9}-27t^{8}+36t^{7}-81t^{6}+36t^{5}-27t^{4}+28t^{3}+12t^{2}+3t+1)^{2}.$
Theorem 5.1.4.
The generating function $f(t)$ for the number of spanning trees in $\mathcal{G}_{5}$ is
$\frac{N_{5}}{D_{5}}$
where $N_{5}$ is a polynomial of degree $161$ and
$D_{5}$ $=(t-1)^{2}(t^{8}+3t^{7}+6t^{6}+10t^{5}+15t^{4}+10t^{3}+6t^{2}+3t+1)^{2}(t^{8}+3t^{7}+6t^{6}-t^{5}+15t^{4}-t^{3}+6t^{2}+3t+1)^{2}$
$(t^{16}-5t^{15}+10t^{14}-10t^{13}-28t^{12}+10t^{11}+110t^{10}+110t^{9}+88t^{8}+110t^{7}+110t^{6}+10t^{5}-28t^{4}-10t^{3}+10t^{2}$
$-5t+1)^{2}(t^{16}-5t^{15}-23t^{14}-10t^{13}-94t^{12}-485t^{11}+242t^{10}+110t^{9}+649t^{8}+110t^{7}+242t^{6}-485t^{5}-94t^{4}$
$-10t^{3}-23t^{2}-5t+1)^{2}(t^{32}+t^{31}+12t^{30}+45t^{29}+45t^{28}-1561t^{27}+3917t^{26}-3222t^{25}-3981t^{24}+7745t^{23}$
$+26379t^{22}-88937t^{21}+84093t^{20}+63864t^{19}-153881t^{18}-202281t^{17}+550163t^{16}-202281t^{15}-153881t^{14}$
$+63864t^{13}+84093t^{12}-88937t^{11}+26379t^{10}+7745t^{9}-3981t^{8}-3222t^{7}+3917t^{6}-1561t^{5}+45t^{4}+45t^{3}$
$+12t^{2}+t+1)^{2}.$

### v3

@@ Abstract
Using the theoretical basis developed by Yao and Zeilberger, we consider certain graph families whose structure results in a rational generating function for sequences related to spanning tree enumeration. Said families are powers of cycles and powers of paths; later, we briefly discuss toru graphs and grid graphs. In each case we know, a priori, that the set of spanning trees of the family of graphs can be described in terms of a finite-state-machine, and hence there is a finite transfer-matrix that guarantees the generating function is rational. Finding this “grammar”, and hence the transfer-matrix is very tedious, so a much more efficient approach is to use experimental mathematics. Since computing numerical determinants is so fast, one can use the matrix tree theorem to generate sufficiently many terms, then fit the data to a rational function. The whole procedure can be done rigorously a posteriori.
@@
1 Introduction
A subgraph $T$ of a graph $G$ such that $T$ is a tree with $V(T)=V(G)$ is called a spanning tree of $G$. If $G$ is connected, simple, and not a tree, then it contains several spanning trees. When we count the spanning trees of a graph in this paper, we consider two isomorphic trees with different vertex labels to be different trees; for example, the complete graph on three vertices $K_{3}$ has three spanning trees, not one. For a connected graph $G$, we denote its number of spanning trees by $\tau(G)$. Cayley’s formula counts the number of spanning trees of a complete graph on $n$ vertices to be $n^{n-2}=\tau(K_{n})$.
The generating function $f(x)$ for a sequence $(a_{0},a_{1},\dots)$ is the formal power series obtained by setting the sequence terms as its coefficients:
@@ 1 Introduction
Whenever a sequence has a recurrence of finite length, with constant coefficients, it is called C-finite. In other words, $(a_{0},a_{1},\dots)$ is C-finite (of order $r$) if there is a fixed $r$ and coefficients $c_{0},\dots,c_{r-1}$ with $c_{0}\neq 0$ such that
$a_{n+r}=c_{r-1}a_{n+r-1}+\dots+c_{0}a_{n}$
@@
[4, Thm 4.3] A sequence $(a_{0},a_{1},\dots)$ is C-finite (of order $r$) with recurrence
$a_{n+r}=c_{r-1}a_{n+r-1}+\dots+c_{0}a_{n}$
@@
$\sum a_{n}x^{n}=\frac{p(x)}{1-c_{r-1}x+\dots-c_{1}x^{r-1}-c_{0}x^{r}}$
for some polynomial $p(x)$ of degree at most r1r-1.
In fact, the polynomial $p(x)$ in the previous theorem is determined by the initial values $(a_{0},\dots,a_{r-1})$ of the sequence. Later, we will introduce various (ordered) families of graphs. Each graph family will give rise to several numerical sequences (such as number of spanning trees) which will be indexed, in some way, by the number of vertices of the underlying graph in the family. To disambiguate the numerator $p(x)$, we will specify the smallest member in each graph family (which corresponds to the first term, $a_{0}$, in any resulting numerical sequence).
Due to the recursive nature of the graph families we consider, we presume that there is a finite transfer matrix which describes the corresponding sequence. (In a paper by Yao and Zeilberger [8], there is an example of such a construction for the family of $k\times n$ grid graphs, for fixed $k$.) Thanks to this, we can find the sequence’s rational generating function by computing a large number of terms in the sequence. The guessing is done in a Maple procedure, described in the same paper [8], called GuessRec.
To generate the terms of whichever sequence we consider, we apply Kirchhoff’s matrix-tree theorem, which allows us to compute the number of spanning trees of a(n $n$-vertex) graph by starting with its Laplacian matrix, taking an n1n-1 by n1n-1 (matrix) minor, and computing its determinant. Section˜3 provides more detail on how to count leaves by applying matrix-tree theorem. Below, we define the Laplacian matrix of a graph for the reader’s convenience.
If $G$ is a (simple) graph with vertices $v_{1},\dots,v_{n}$, its Laplacian matrix is a symmetric matrix with its entries defined by
@@
If the reader wishes to further review labelled spanning trees, we direct them to classics such as [1] or [6].
In Section˜3, we describe methods to count the leaves (across all spanning trees) for graphs in the same families. For Section˜4, we consider an asymptotic constant relating total number of leaves to total number of spanning trees in any families (for which the constant is well-defined). We state our main experimental results in Section˜5 and Section˜6, then briefly consider an additional pair of graph families in Section˜7. In Section˜8, we describe how to numerically estimate the aforementioned constant with random sampling and an implementation to make such estimates. We developed a Maple package that carries out the computations in this paper. It is described in Section˜9 and available at [2]. Finally, we conclude with some conjectures in Section˜10.
@@ 2 Preliminaries
For ease of notation, we assume all graphs are simple. We write $n]:=\{1,\dots,n\}$.
Definition. For a graph $G$, the distance $|u,v|$ between two vertices $u,v$ is the length of the shortest path between them. If $u$ and $v$ are in different components of $G$, we say the distance between them is infinite and write $|u,v|=\infty$.
Definition. Let $G$ be a graph. For an integer $k\geq 1$, the $k$-th power of $G$, denoted $G^{k}$, is the graph obtained from $G$ such that $V(G^{k}):=V(G)$ and $E(G^{k}):=\{uv:u,v\in V(G),\ 1\leq|u,v|\leq k\}$. See Figure˜1.
From the definition, we see that if $d$ is the diameter of a connected graph $G$ (defined as $d:=\max_{u,v\in V(G)}|u,v|$ and denoted by $\text{diam }G$), then $G^{d}$ is the complete graph on $|V(G)|$ vertices. In some cases, it’s natural to exclude complete graphs and hence consider only $G^{k}$ when $k<d$. Additionally, one may note that if there are two shortest paths (in $G$) between two vertices $u,v$ of length $m>1$, then $G^{m}$ may be better viewed as a multigraph where there are two edges between $u,v$.
@@
The condition in each construction ensures that our graph families do not contain any “unnatural" complete graphs. For those cases, we can once again recall Cayley’s formula: $\tau(K_{n})=n^{n-2}$. Since the diameter of a cycle (resp. path) increases with the number of vertices, the condition $k<\text{diam }C_{n}=\lfloor n/2\rfloor$ is implicitly a lower bound on the number of vertices. Explicitly,
$\mathcal{C}_{k}$ $=\{{C_{n}}^{k}\ :\ n\geq 2k+1\}$
$\mathcal{P}_{k}$ $=\{{P_{n}}^{k}\ :\ n\geq k+2\}.$
As discussed in the introduction, the initial values of a C-finite sequence determine the numerator of its rational generating function. The minimum values of $n$ above will mark the beginning of that family’s sequence. Since paths and cycles are ubiquitous in graph theory, they are the focus of this paper. We provide a quick application below.
Let $G$ be a connected graph. Suppose that we wish to estimate the number of spanning trees of its $k$-th power graph, $\tau(G^{k})$. One way to estimate $\tau(G^{k})$ is to fix a vertex $v\in V(G)$, let vertex sets $H_{1},\dots,H_{N}$ be the components of GvG-v and $m_{i}(v):=\max_{u\in H_{i}}|u,v|$. Then, deduce the bound:
@@ 3 Counting Total Number of Leaves
The next proposition plays a key role in our implementation for computing a parameter for graph families that we call the B-Z constant. The idea is to count the leaves (across all spanning trees) in a graph by counting how many times each vertex appears as a leaf. For each vertex in $G$, remove the vertex from $G$ and count the spanning trees in the resulting graph. (This method appears in [1].)
@@
∎ We say $G$ is vertex-transitive if for any $u,v\in V(G)$ there is an automorphism $\varphi$ of $G$ such that $\varphi(u)=v$ and $\varphi(v)=u$.
@@ 4 The B-Z Constant: Average Number of Leaves per Vertex
whenever the limit exists. This is the limit of the average number of leaves in a random spanning tree (and normalized by number of vertices) in the discussed family. For the infinite family of complete graphs $K_{n}$, the limit is famously equal to $1/e$ (see below). In general, we have the bound $\frac{\sum|\mathcal{L}(T)|}{n\tau(G)}\leq 1$ for any $G$. The rest of the section provides a few examples of B-Z constants and our approach for computing them.
The B-Z constants for both the family of cycles and the family of paths is equal to 0, since every tree (a path) has a constant number of leaves (two). A star graph is a connected graph where all edges share the same vertex. When $\mathcal{G}$ is the family of star graphs, its B-Z constant is $1$ because the unique tree of a star graph (itself) has n1n-1 leaves.
As described in Section˜3, it is possible to compute $\sum|\mathcal{L}(T)|$ by applying matrix-tree theorem. In the case of vertex-transitive graphs, computation of $\sum|\mathcal{L}(T)|$ (and hence the B-Z constant) is made faster thanks to Corollary˜3.0.2. We used this fact in the cases of: powers of cycles ($\mathcal{P}_{k}$) and torus graphs.
Incidentally, it’s easy to compute the B-Z constant for complete graphs by using Corollary˜3.0.2 in conjunction with Cayley’s formula:
@@
Recall Cayley’s formula, which states $\tau(K_{n})=n^{n-2}$ for $n\geq 2$. With $\mathcal{T}:=\mathcal{T}(K_{n})$, Corollary˜3.0.2 tells us that $\sum|\mathcal{L}(T)|=n\cdot(n-1)^{n-2}$. Hence,
@@
From a theorem in a joint paper by Hladký, Nachmias, and Tran [3, Thm. 1.5], a stronger result follows. The B-Z constant for a family of graphs, where each member has minimum degree $\delta n$ (for a fixed $\delta>0$), is $e^{-1}$. Their paper uses probabilistic methods, as well as graphons in the setting of dense graphs, to obtain this result. Each family $\mathcal{F}$ of graphs we considered in this paper is much sparser: there is a constant $c$ for which all $G\in\mathcal{F}$ satisfy $\Delta(G)\leq c$; in other words, the maximum degree sequence $\{\Delta(G_{n})\ :G_{n}\in\mathcal{F}\}$ is bounded by a constant. In the case of $\mathcal{G}_{k}$ and $\mathcal{P}_{k}$, a (tight) bound is $c=2k$; whereas $c=4$ is tight for families of torus graphs and grid graphs.
Earlier, we noted that star graphs have B-Z constant equal to $1$. The next proposition observes that: for any rational number (in the interval $[0,1]$), we can find a graph family with B-Z constant equal to it. Such a family is constructed below by slightly modifying the family of star graphs.
@@
To construct each member $G_{n}\in\mathcal{G}$ (assuming $n\geq 1$): begin with a star graph $S_{pn}$ with $pn$ leaves; then, subdivide (qp)n(q-p)n edges in $S_{pn}$ and call this $G_{n}$. Note that $|V(G_{n})|=qn+1$ and that $G_{n}$ has $pn$ leaves. Furthermore, $G_{n}$ is a tree so it has exactly one spanning tree. It follows that the B-Z constant for $\mathcal{G}$ is $\lim_{n\to\infty}\frac{pn}{qn+1}=\frac{p}{q}$.∎
Unlike these modified star graph families, the families we consider in upcoming sections have B-Z constants that are algebraic, but not rational.
@@ 5 Powers of a Cycle: Experimental Results
In upcoming sections, we determine the B-Z constants of several graph families. It’s possible, as we do in later sections, to compute the B-Z constant of a graph family from its asymptotic behavior in the total number of leaves and number of spanning trees (which can be obtained from their respective generating functions). We used standard residue calculations, implemented in procedure BZc from our Maple package (Section˜9).
@@ 5.1 Generating Functions for the Number of Spanning Trees
In upcoming sections, we determine the B-Z constants of several graph families. It’s possible, as we do in later sections, to compute the B-Z constant of a graph family from its asymptotic behavior in the total number of leaves and number of spanning trees (which can be obtained from their respective generating functions). We used standard residue calculations, implemented in procedure BZc from our Maple package (Section˜9).